An investor short sells 400 shares of a stock for $ 20.62 per share. The initial margin is 50 %​, and the maintenance margin is 29 %. The price of the stock rises to $ 29.95 per share. What is the​ margin, and will there be a margin​ call?

The margin in the account is _______________%. ​ (Round to the nearest​ percent.)
Answer #1

Initial margin = 400*20.62*50% =$4124

Loss =(20.62-29.95)*400 =$3732

Balance in margin account =$392

I.e. 392/400*20.62 = 4.75%

I.e. 5%

Yes there will be margin call since margin has fallen below maintenance margin level